Your current income is​ $50,000 per​ year, and you would like to maintain your current standard of living​ (i.e., your purchasing​ power) when you retire. If you expect to retire in 30 years and expect inflation to average​ 3% over the next 30​ years, what amount of annual income will you need to live at the same comfort level in 30​ years?
Answer #2
Answer #1

FV = PV (1+r)t

Where, PV = Present value, r = rate of interest, t = number of years.

FV = 50000 (1+.03)30

FV = 50000 * 2.427262

FV = $121363

I will need $121363 after 30 years that will be same comfort level.